-7m-2p=-13
m-2p=11

Answer by diazabdulm(7)

So far, it looks like the best way to solve this system is by elimination of the 2p
Multiply either equation by -1, in this case I choose the first equation
-1(-7m-2p=-13)
m-2p=11
that would equal
7m+2p=13
m-2p=11
Add the two equations and get
8m=24
Divide by 8
m=3

Now after I know one number, I substitute that number into either equation to know the second number, I choose the second equation

m=3
m-2p=11
Substitue m=3
3-2p=11
Subtract 3
-2p=8
Divide by -2

p=-4
So
m=3 and p=-4